5. Given sec θ = 13/12 Calculate all other trigonometric ratios

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
We know that sec function is the reciprocal of the cos function which is equal to the ratio of the length of the hypotenuse side to the adjacent side Let us assume a right angled triangle ABC, right angled at B sec θ =13/12 = Hypotenuse/Adjacent side = AC/AB Let AC be 13k and AB will be 12k Where, k is a positive real number. According to the Pythagoras theorem, the squares of the hypotenuse side is equal to the sum of the squares of the other two sides of a right angle triangle and we get, AC2=AB2 + BC2 Substitute the value of AB and AC (13k)2= (12k)2 + BC2 169k2= 144k2 + BC2 169k2= 144k2 + BC2 BC2 = 169k2 – 144k2 BC2= 25k2 Therefore, BC = 5k Now, substitute the corresponding values in all other trigonometric ratios So, Sin θ = Opposite Side/Hypotenuse = BC/AC = 5/13 Cos θ = Adjacent Side/Hypotenuse = AB/AC = 12/13 tan θ = Opposite Side/Adjacent Side = BC/AB = 5/12 Cosec θ = Hypotenuse/Opposite Side = AC/BC = 13/5 cot θ = Adjacent Side/Opposite Side = AB/BC = 12/5